The height of transmission tower at anyvplace on the earth 245m the maximum distance up to which transmission of tower will reach is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Let the Radius of the earth be R.
then, the relation b/w the R,h and the maximum distance up to which transmission of tower will reach is:
245^2 + (6400*1000)^2 = x^2
we get,
x^2 = 60025 + 4096 x 10^10
x = root( 60025 + 4096 x 10^10) meter. is the required distance.